OpenStudy (yinshy):

For the investment with interest compounded annually, find the final balance and the total interest earned. $500 for 2 years at 8% A. final balance: $540 interest earned: $40 B. final balance: $580 interest earned: $80 C. final balance: $583.20 interest earned: $83.20 D. final balance: $620 interest earned: $120

OpenStudy (mathewr.):

B

OpenStudy (yinshy):

may i ask how?

OpenStudy (mathewr.):

It has an 8% interest so you multiply 500 x .08 =40 so its 40 for 1 year x 2 = 80

OpenStudy (mathstudent55):

Answer is not B. The interest rate is 8%, so after 1 year, multiply the principal by 1.08. Then take the new amount, and multiply by 1.08 again to find the amount at the end of the second year.

OpenStudy (yinshy):

C?

OpenStudy (mathstudent55):

@MathewR. That is simple interest. The problem here states compound interest, so you earn interest on the interest.

OpenStudy (mathstudent55):

500 * 1.08 * 1.08 = 583.20

OpenStudy (mathstudent55):

C is correct.

OpenStudy (mathewr.):

ahh i didn't read the compound part... My bad i apologize!

OpenStudy (yinshy):

its cool

OpenStudy (mathstudent55):

Start: $500 After 1 year: $500 * 1.08 = $540 Now start at the end of 1 year: $540 After the second year: $540 * 1.08 = $583.20